OpenStudy (anonymous):

Which of the following shows a proportion? four over nine equals eighteen over eight three over eight equals six over four six over twelve equals eight over four nine over fifteen equals twelve over twenty

OpenStudy (anonymous):

plz help

OpenStudy (luigi0210):

Try writing them out?

OpenStudy (anonymous):

4/9=18/8 3/8=6/4 6/12=8/4 9/15=12/20

OpenStudy (anonymous):

wont anyone help this poor lost soul

ganeshie8 (ganeshie8):

cross multiply and see

OpenStudy (anonymous):

im not good at that

ganeshie8 (ganeshie8):